# for Generative QA

### v1

@@ Abstract
To use generative question-and-answering (QA) systems for decision-making and in any critical application, these systems need to provide well-calibrated confidence scores that reflect the correctness of their answers. Existing calibration methods aim to ensure that the confidence score is on average indicative of the likelihood that the answer is correct. We argue, however, that this standard (average-case) notion of calibration is difficult to interpret for decision-making in generative QA. To address this, we generalize the standard notion of average calibration and introduce $\beta$-calibration, which ensures calibration holds across different question-and-answer groups. We then propose discretized posthoc calibration schemes for achieving $\beta$-calibration.
Keywords: calibration; question-answering
@@ 1 Introduction
To evaluate whether the obtained confidence scores are actually well-calibrated, a common criterion is the expected (average-case) calibration error (Tian et al., 2023; Xiong et al., 2024). Suppose a model claims that its answer has a confidence of $p$. Based on only this one answer it is not possible to know whether this confidence was well-calibrated or not. But when considering multiple question-answer pairs, let’s say $N_{p}$ with claimed confidence $p$, we can verify how many answers were actually correct and measure the error between the claimed confidence and the model’s accuracy. By averaging over all these errors, we can measure the calibration of the model’s confidence scores. [^1] Note that the accuracy of the LM is itself unaffected by calibration, as the latter does not change the weights of the LM model.
While this average-case calibration measure makes sense for models trained and evaluated on specific tasks, its applicability for generative QA is highly questionable due to the averaging is now over all QA pairs. The reason being that generative QA systems can be applied in various domains and topics, e.g., to answer questions about geography as well as about politics or medicine. Consider, for instance, the QA pairs shown in Figure 1. On average this model has a calibration error of $0.5$. But as far as User 1 is concerned, the calibration is much worse with an error of $0.8$. User 2, on the other hand, makes a completely different experience, as for them the model seems to be much better calibrated than indicated by the average calibration error. This motivates our notion of $\beta$-calibration, where the calibration target is conditional on the group of the question-and-answer pair.
@@
A limitation of relying on LM-elicited confidence directly, or post-hoc calibrated using temperature scaling  (Tian et al., 2023), is that the output probability is not discretized, making performance difficult to assess  (Kumar et al., 2019). We overcome this problem by developing posthoc $\beta$-calibration schemes on elicited confidence scores that use ideas of (histogram) binning (Zadrozny and Elkan, 2001) and scaling-binning (Kumar et al., 2019) that are by construction discretized.
@@
    We define $\beta$-calibration, a principled and interpretable notion of calibration in the generative QA setting. The $\beta$ here refers to any fixed mapping of all possible question-and-answer pairs to a finite set. Our $\beta$-calibration notion generalizes the standard average-case calibration notion by requiring calibration conditional on $\beta$. Due to this conditioning on $\beta$, the guarantee of $\beta$-calibration is stronger than the standard calibration guarantee. By instantiating this framework with different $\beta$’s, we have the flexibility of defining question-and-answer groups across which we would like calibration guarantees to hold. In this paper, we present an instantiation of $\beta$ as a kd-tree, an adaptive multivariate histogram method that provides cohesive groupings of QA pairs.
@@
    We propose two posthoc calibration techniques for $\beta$-calibration: (a) $\beta$-binning and (b) scaling-$\beta$-binning. The latter uses the former as a subroutine. This is particularly useful in the practical setting where some $\beta$-induced groups may lack data. We show that both methods satisfy a distribution-free approximate $\beta$-calibration guarantee. In both cases, the approximation level is used to decide on the number of points per bin, a key hyperparameter.
@@
    Finally, we experiment on newly developed elicitation prompts across 5 commonly used QA benchmark datasets of various sizes. We find that $\beta$-binning and scaling-$\beta$-binning achieve lower $\beta$-calibration error than elicited confidence scores and other baselines. We further justify our framework by demonstrating its performance in a downstream selective QA task. Our posthoc calibration algorithms achieve around 10-40% increase in $\beta$-calibration performance and up to 30% increase in selective answering performance.
2 Defining $\beta$-Calibration
@@ 2.1 Notation
$P:=P_{Q}\times P_{A|q}\times P_{H|q,a}\times P_{Y|q,a}\leavevmode\nobreak\ .$ (1)
@@
In words, that the conditional distribution on $Y$ conditional on the prediction that $h(Q,A)=p$ is a Bernoulli distribution with bias $p$. In the following, we refer to Eq. ( 2) as average-case calibration to distinguish it from $\beta$-calibration.
@@
2.2 $\beta$-calibration Instead of Standard Calibration
@@
Definition 2.3.
($\beta$-calibration) $h$ is $\beta$-calibrated for the distribution $P$ in Eq. ( 1) if:
@@
a.s. for all $p\in[0,1]$, where $\beta:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow{\mathcal{S}}$ for some embedding space ${\mathcal{S}}$.
In this work, we focus on a finite set ${\mathcal{S}}$ and $\beta$ is a (deterministic) discretization scheme that maps $(q,a)$ to a finite set $S$, i.e., $\beta$ induces a partitioning of the QA space according to the output of $\beta$. Note that this $\beta$-calibration reduces to calibration in Eq. ( 2) if $\beta(q,a)$ is the same for all question-and-answers $(q,a)$. Intuitively, $\beta$ is chosen such that the pre-image of a specific value of $\beta$ represents a grouping that an end-user might be interested in.
Going back to Example 2.1, $\beta$-calibration would tell us the following: “among question and answer pairs with confidence $p$ that are mapped to the same $\beta$ value (as the user’s $(q,a)$ pair), the fraction who are correct ($y=1$) is also $p$”. Further, we also extend Eq. ( 3) to propose an error metric for $\beta$-calibration:
Definition 2.4 ($\beta$-calibration error).
The $\beta$-calibration error of $h$ is defined as:
@@
2.3 Generalizing (Average) Calibration via kd-Tree Instantiation of $\beta$
The definition of $\beta$-calibration (Definition 2.3), requires a predefined $\beta$ mapping. While our schemes developed in Section 3 can work with any arbitrary $\beta$ mapping, we propose a general approach for $\beta$ in the language model QA setting: the embed-then-bin method. This involves computing vector embeddings from question-and-answer pairs and then creating a multidimensional histogram over these embeddings. The range of $\beta$ is then set to index over the histogram bins of the embeddings. We write $\beta=\beta_{\text{hist}}\circ\beta_{\text{emb}}$, where $\beta_{\text{emb}}:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow\mathbb{R}^{M}$ computes an $M$-dimensional embedding and $\beta_{\text{hist}}:{\mathbb{R}}^{M}\rightarrow{\mathcal{S}}$ computes the index of the embedding histogram bin containing the embedding.
Specifically, we use the [CLS] token embedding from the pre-trained DistilBERT model (Sanh et al., 2019): $\beta_{\text{emb}}:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow\mathbb{R}^{768}$ ($M=768$). As for the histogram, we adapt a kd-tree (Bentley, 1975) to bin each vector to an integer representing the index of the partition containing the vector: $\beta_{\text{hist}}:{\mathbb{R}}^{768}\rightarrow S$, where $S$ is the set of partition indices. See Appendix C for details on our kd-tree construction. An important hyperparameter is the maximum depth of the tree $d$, which determines the number of partitions.
@@ 2.3 Generalizing (Average) Calibration via kd-Tree Instantiation of $\beta$
The choice of kd-tree generalizes the standard calibration as when $d=0$, $\text{CE}(h;\beta)$ reduces to $\text{CE}(h)$. The hyperparameter $d$ should be chosen based on the downstream metric that needs to be optimized. In the experiments section, we further describe how $\beta_{\text{hist}}$ are constructed.
3 Achieving Posthoc $\beta$-Calibration
@@
As discussed earlier, an elicited confidence score function ($h$) of LM-model is not guaranteed to be $\beta$-calibrated (or even calibrated under average-case). For any $\beta:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow{\mathcal{S}}$, the goal of posthoc $\beta$-calibration is to design a scheme that that can take any $h$ and transform it to a $\beta$-calibrated confidence score. We wish to learn a posthoc calibrator $g:[0,1]\rightarrow[0,1]$ such that $g\circ h$ is (approximately) $\beta$-calibrated using a calibration dataset $D=\{(q_{i},a_{i},h_{i},y_{i})\}_{i\in[N]}$. Additional missing details from this section are collected in Appendix A.
@@
Our recalibration schemes utilize the existing building blocks in posthoc calibration literature, like histogram binning (Zadrozny and Elkan, 2001) and scaling (Platt, 1999). The novelty of our recalibration methods lies how we generalize and combine these building blocks achieving the best of both worlds: our binning component generalizes histogram binning to any partitioning and enables distribution-free guarantees, and our hierarchical scaling component reduces overfitting, enabling learning to be performed across partitions which may have varied number of data points.
@@
Our schemes take the $\beta$ function as input, and all our schemes and theoretical results hold for any $\beta$. In our experiments, we will use the kd-trees based $\beta$ instantiation from Section 2.3. We focus on methods that output discretized scores, since this has been shown to be easier to assess (Appendix A.1). We achieve this by first discretizing the QA space using $\beta$ and then ensuring that for each QA partition $s\in{\mathcal{S}}$ we output a calibrated confidence score. We start by describing an adaptation of the classical histogram binning (Zadrozny and Elkan, 2001) idea for achieving $\beta$-calibration. In Section 3.2, we build upon this to provide a more robust algorithm that also uses S before binning.
3.1 Approach 1: $\beta$-binning
@@
$\beta$-binning is a standalone posthoc $\beta$-calibration algorithm. It uses as a subroutine, uniform-mass-double-dipping histogram binning(UMD) (Gupta and Ramdas, 2021), which we describe in Algorithm 4. Informally, the UMD procedure partitions the interval $[0,1]$ into $B$ bins using the histogram of $h$ values from the calibration dataset $D$, ensuring that each bin has approximately the same number of calibration points. It returns a calibrator function $g_{\text{UMD}}$ that takes as input an uncalibrated confidence score, then allocates it to one of the $B$ bins, and returns the probability of label being $1$ (estimated as the average of the $y$ values from the calibration dataset $D$ that are mapped into that bin).
@@
In Algorithm 1, we present a subroutine that uses the UMD procedure to construct a different calibrator for each QA-partition, which we will invoke with different inputs. Algorithm 1 takes as input a dataset ${\mathcal{D}}=\{\left(q_{i},a_{i},h_{i},t_{i}\right)\}_{i\in[n]}$ where $h_{i}$ is the (elicited) confidence score for the answer $a_{i}$ and $t_{i}$ is the (potentially noisy) label of the correctness of answer $a_{i}$ for the question $q_{i}$. Algorithm 1 partitions the input dataset ${\mathcal{D}}$ to $\{{\mathcal{D}}_{s}\}_{s\in{\mathcal{S}}}$, where ${\mathcal{D}}_{s}=\{\left(q_{i},a_{i},h_{i},t_{i}\right):\beta(q_{i},a_{i})=s\}$. For each $s\in{\mathcal{S}}$, UMD is fit using ${\mathcal{D}}_{s}$ to construct a $g_{\text{UMD}}$ calibrator that is stored in a hash table ${\mathcal{G}}$, keyed by $s$. We next describe the training and test processes of $\beta$-binning.
Algorithm 1 $\beta$-binning: Train-time Subroutine
@@
Algorithm 2 $\beta$-binning: Test-time Subroutine
@@
0:  (Approximately) $\beta$-calibrated confidence score
@@
Test Process. For a test input $(q_{\text{test}},a_{\text{test}},h_{\text{test}})$, we invoke Algorithm 2. A fitted UMD ($g_{\text{UMD}}$) for this $(q_{\text{test}},a_{\text{test}})$ pair is retrieved from the hash table ${\mathcal{G}}$ using $\beta(q,a)$ as key, which is then invoked to obtain a $\beta$-calibrated confidence score $g_{\text{UMD}}(h_{\text{test}})$. In our kd-tree instantation, if ${\mathcal{G}}$ does not have a calibrator for a $\beta(q_{\text{test}},a_{\text{test}})$, we use UMD to calibrate that pair. This corresponds to points that lie outside of the bounded kd-tree spaces (refer to Appendix C). Since we can generate a large amount of question-and-answer pairs without requiring ground truth, the number of test points in our experiments that fall outside the bounded spaces is negligible.
@@
Note that while we assumed here access to the true labels $y_{i}$’s, we can also operate $\beta$-binning with proxy labels generated by say another LM on QA pairs, as discussed more in following subsections. In practice, users are likely to define large $S$ (for example, large maximum depth of tree $d$) in order to obtain more cohesive groups of QA pairs. UMD, and therefore $\beta$-binning, may overfit if the number of data points within ${\mathcal{D}}_{s}$ is small — a highly likely scenario if $\beta$ induces very fine partitions. We overcome these issues with our next approach that combines scaling with binning.
3.2 Approach 2: Scaling-$\beta$-binning

### v2

# QA-Calibration of Language Model Confidence Scores

@@ Abstract
To use generative question-and-answering (QA) systems for decision-making and in any critical application, these systems need to provide well-calibrated confidence scores that reflect the correctness of their answers. Existing calibration methods aim to ensure that the confidence score is on average indicative of the likelihood that the answer is correct. We argue, however, that this standard (average-case) notion of calibration is difficult to interpret for decision-making in generative QA. To address this, we generalize the standard notion of average calibration and introduce QA-calibration, which ensures calibration holds across different question-and-answer groups. We then propose discretized posthoc calibration schemes for achieving QA-calibration. We establish distribution-free guarantees on the performance of this method and validate our method on confidence scores returned by elicitation prompts across multiple QA benchmarks and large language models (LLMs).
@@ 1 Introduction
To evaluate whether the obtained confidence scores are actually well-calibrated, a common criterion is the expected (average-case) calibration error (Tian et al., 2023; Xiong et al., 2024). Suppose a model claims that its answer has a confidence of $p$. Based on only this one answer, it is not possible to know whether this confidence was well-calibrated or not. But when considering multiple question-answer pairs–say $N_{p}$ pairs each with a claimed confidence $p$, we can verify how many answers are actually correct and measure the error between the claimed confidence and the model’s accuracy. Averaging over these errors quantifies of the model’s confidence scores. [^1] Note that the accuracy of the LM is itself unaffected by calibration, as calibration does not change the weights of the LM model.
While this average-case calibration measure makes sense for models trained and evaluated on specific tasks, its applicability to generative QA is questionable because it averages over all QA pairs. This is because generative QA systems can be applied across various domains and topics, such as geography, politics, or medicine. Consider, for instance, the QA pairs shown in Figure 1. On average, this model has a calibration error of $0.5$. From User 1’s perspective, the calibration is much worse, with an error of 0.8. Meanwhile, User 2 has a completely different experience, finding the model to be much better calibrated than indicated by the average calibration error. This motivates our notion of QA-calibration, where the calibration target is conditional on the group of the question-and-answer pair.
@@
A limitation of relying on LM-elicited confidence directly, or posthoc calibrated using temperature scaling  (Tian et al., 2023), is that the output probability is not discretized, making performance difficult to assess  (Kumar et al., 2019). We overcome this problem by developing posthoc QA-calibration schemes on elicited confidence scores that use ideas of (histogram) binning (Zadrozny & Elkan, 2001) and scaling-binning (Kumar et al., 2019) that are by construction discretized.
@@
    We introduce QA-calibration, a principled and interpretable notion of calibration in generative QA. Crucially, it defines a fixed mapping $\beta$ from all possible question–and–answer pairs to a finite set. Our QA-calibration notion generalizes the standard average-case calibration notion by requiring calibration conditional on $\beta$. Due to this conditioning on $\beta$, the guarantee of QA-calibration is stronger than the standard calibration guarantee. By instantiating this framework with different $\beta$’s, we have the flexibility of defining question-and-answer groups across which we would like calibration guarantees to hold. In this paper, we present an instantiation of $\beta$ as a kd-tree, an adaptive multivariate histogram method that provides cohesive groupings of QA pairs.
@@
    We propose two posthoc calibration techniques for QA-calibration: (a) QA binning and (b) scaling QA binning. The latter uses the former as a subroutine. This is particularly useful in the practical setting where some QA induced groups may lack data. We show that both methods satisfy a distribution-free approximate QA-calibration guarantee. In both cases, the approximation level is used to decide on the number of points per bin, a key hyperparameter.
@@
    Finally, we experiment on newly developed elicitation prompts across 5 commonly used QA benchmark datasets of various sizes. We find that QA binning and scaling QA binning achieve lower QA-calibration error than elicited confidence scores and other baselines. We further justify our framework by demonstrating its performance in a downstream selective QA task. Our posthoc calibration algorithms achieve around 10-40% increase in QA-calibration performance and up to 30% increase in selective answering performance.
2 Defining QA-calibration
@@ 2.1 Notation
$P:=P_{Q}\times P_{A|q}\times P_{H|q,a}\times P_{Y|q,a}~{}.$ (1)
@@
In words, that the conditional distribution on $Y$ conditional on the prediction that $h(Q,A)=p$ is a Bernoulli distribution with bias $p$. In the following, we refer to Eq. ( 2) as average-case calibration to distinguish it from QA-calibration.
@@
2.2 QA-calibration Instead of Standard Calibration
@@
Definition 2.3 (QA-calibration).
$h$ is QA calibrated for the distribution $P$ in Eq. ( 1) if:
@@
a.s. for all $p\in[0,1]$, where $\beta:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow{\mathcal{S}}$ is a fixed mapping to some embedding space ${\mathcal{S}}$.
In this work, we focus on a finite set ${\mathcal{S}}$ and $\beta$ is a (deterministic) discretization scheme that maps $(q,a)$ to a finite set $S$, i.e., $\beta$ induces a partitioning of the QA space according to the output of $\beta$. Note that this QA-calibration reduces to calibration in Eq. ( 2) if $\beta(q,a)$ is the same for all question-and-answers $(q,a)$. Intuitively, $\beta$ is chosen such that the pre-image of a specific value of $\beta$ represents a grouping that an end-user might be interested in.
Going back to Example 2.1, QA-calibration would tell us the following: “among question and answer pairs with confidence $p$ that are mapped to the same $\beta$ value (as the user’s $(q,a)$ pair), the fraction who are correct ($y=1$) is also $p$”. Further, we also extend Eq. ( 3) to propose an error metric for QA-calibration:
Definition 2.4 (QA-calibration error).
The QA-calibration error of $h$ with fixed mapping $\beta$ is defined as:
@@
2.3 Generalizing (Average-case) Calibration via kd-tree Instantiation of $\beta$
The definition of QA-calibration (Definition 2.3), requires a predefined $\beta$ mapping. While our schemes developed in Section 3 can work with any arbitrary $\beta$ mapping, we propose a general approach for $\beta$ in the language model QA setting: the embed-then-bin method. This involves computing vector embeddings from question-and-answer pairs and then creating a multidimensional histogram over these embeddings. The range of $\beta$ is then set to index over the histogram bins of the embeddings. We write $\beta=\beta_{\text{bin}}\circ\beta_{\text{emb}}$, where $\beta_{\text{emb}}:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow\mathbb{R}^{M}$ computes an $M$-dimensional embedding and $\beta_{\text{bin}}:{\mathbb{R}}^{M}\rightarrow{\mathcal{S}}$ computes the index of the embedding histogram bin containing the embedding.
Specifically, we use the [CLS] token embedding from the pre-trained DistilBERT model (Sanh et al., 2019): $\beta_{\text{emb}}:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow\mathbb{R}^{768}$ ($M=768$). As for the histogram, we adapt a kd-tree (Bentley, 1975) to bin each vector to an integer representing the index of the partition containing the vector: $\beta_{\text{bin}}:{\mathbb{R}}^{768}\rightarrow S$, where $S$ is the set of partition indices. See Appendix C for details on our kd-tree construction. An important hyperparameter is the maximum depth of the tree $d$, which determines the number of partitions.
@@ 2.3 Generalizing (Average-case) Calibration via kd-tree Instantiation of $\beta$
The choice of kd-tree generalizes the standard calibration as when $d=0$, $\text{CE}(h;\beta)$ reduces to $\text{CE}(h)$. The hyperparameter $d$ should be chosen based on the downstream metric that needs to be optimized. In the experiments section, we further describe how $\beta_{\text{bin}}$ are constructed.
3 Achieving Posthoc QA-calibration
@@
As discussed earlier, an elicited confidence score function ($h$) of LM-model is not guaranteed to be QA calibrated (or even calibrated under average-case). For any $\beta:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow{\mathcal{S}}$, the goal of posthoc QA-calibration is to design a scheme that that can take any $h$ and transform it to a QA calibrated confidence score. We wish to learn a posthoc calibrator $g:[0,1]\rightarrow[0,1]$ such that $g\circ h$ is (approximately) QA calibrated using a calibration dataset $D=\{(q_{i},a_{i},h_{i},y_{i})\}_{i\in[N]}$. Additional missing details from this section are collected in Appendix A.
@@
Our recalibration schemes utilize the existing building blocks in posthoc calibration literature, like histogram binning (Zadrozny & Elkan, 2001) and scaling (Platt, 1999). The novelty of our recalibration methods lies how we generalize and combine these building blocks achieving the best of both worlds: our binning component generalizes histogram binning to any partitioning and enables distribution-free guarantees, and our hierarchical scaling component reduces overfitting, enabling learning to be performed across partitions which may have varied number of data points.
@@
Our schemes take the $\beta$ function as input, and all our schemes and theoretical results hold for any $\beta$. In our experiments, we will use the kd-trees based $\beta$ instantiation from Section 2.3. We focus on methods that output discretized scores, since this has been shown to be easier to assess (Appendix A.1). We achieve this by first discretizing the QA space using $\beta$ and then ensuring that for each QA partition $s\in{\mathcal{S}}$ we output a calibrated confidence score. We start by describing an adaptation of the classical histogram binning (Zadrozny & Elkan, 2001) idea for achieving QA-calibration. In Section 3.2, we build upon this to provide a more robust algorithm that also uses scaling before binning.
3.1 Approach 1: QA binning
@@
QA binning is a standalone posthoc QA-calibration algorithm. It uses as a subroutine, uniform-mass-double-dipping histogram binning (UMD) (Gupta & Ramdas, 2021), which we describe in Algorithm 4. Informally, the UMD procedure partitions the interval $[0,1]$ into $B$ bins using the histogram of $h$ values from the calibration dataset $D$, ensuring that each bin has approximately the same number of calibration points. It returns a calibrator function $g_{\text{UMD}}$ that takes as input an uncalibrated confidence score, then allocates it to one of the $B$ bins, and returns the probability of label being $1$ (estimated as the average of the $y$ values from the calibration dataset $D$ that are mapped into that bin).
@@
In Algorithm 1, we present a subroutine that uses the UMD procedure to construct a different calibrator for each QA-partition, which we will invoke with different inputs. Algorithm 1 takes as input a dataset ${\mathcal{D}}=\{\left(q_{i},a_{i},h_{i},t_{i}\right)\}_{i\in[n]}$ where $h_{i}$ is the (elicited) confidence score for the answer $a_{i}$ and $t_{i}$ is the (potentially noisy) label of the correctness of answer $a_{i}$ for the question $q_{i}$. Algorithm 1 partitions the input dataset ${\mathcal{D}}$ to $\{{\mathcal{D}}_{s}\}_{s\in{\mathcal{S}}}$, where ${\mathcal{D}}_{s}=\{\left(q_{i},a_{i},h_{i},t_{i}\right):\beta(q_{i},a_{i})=s\}$. For each $s\in{\mathcal{S}}$, UMD is fit using ${\mathcal{D}}_{s}$ to construct a $g_{\text{UMD}}$ calibrator that is stored in a hash table ${\mathcal{G}}$, keyed by $s$. We next describe the training and test processes of QA binning.
Algorithm 1 QA binning: Train-time Subroutine
@@
Algorithm 2 QA binning: Test-time Subroutine
@@
0:  (Approximately) QA calibrated confidence score
@@
Test Process. For a test input $(q_{\text{test}},a_{\text{test}},h_{\text{test}})$, we invoke Algorithm 2. A fitted UMD ($g_{\text{UMD}}$) for this $(q_{\text{test}},a_{\text{test}})$ pair is retrieved from the hash table ${\mathcal{G}}$ using $\beta(q,a)$ as key, which is then invoked to obtain a QA calibrated confidence score $g_{\text{UMD}}(h_{\text{test}})$. In our kd-tree instantation, if ${\mathcal{G}}$ does not have a calibrator for a $\beta(q_{\text{test}},a_{\text{test}})$, we use UMD to calibrate that pair. This corresponds to points that lie outside of the bounded kd-tree spaces (refer to Appendix C). Since we can generate a large amount of question-and-answer pairs without requiring ground truth, the number of test points in our experiments that fall outside the bounded spaces is negligible.
@@
Note that while we assumed here access to the true labels $y_{i}$’s, we can also operate QA binning with proxy labels generated by say another LM on QA pairs, as discussed more in following subsections. In practice, users are likely to define large $S$ (for example, large maximum depth of tree $d$) in order to obtain more cohesive groups of QA pairs. UMD, and therefore QA binning, may overfit if the number of data points within ${\mathcal{D}}_{s}$ is small — a highly likely scenario if $\beta$ induces very fine partitions. We overcome these issues with our next approach that combines scaling with binning.
3.2 Approach 2: Scaling QA binning